Probability of each question correct is 1/4 or 0.25. Since there are 5 questions, raise 0.25 to the 5th power or (0.25)5. So, probability all correct is 0.0009765.
If you tossed a number cube What is the probability you would toss an 8?
If the "number cube" you are referring to is what we normally call a die (one of a pair of dice), the chances of tossing an 8 are nothing since the cube has only 6 sides and is numbered 1 to 6.
If a "number cube" is something else, perhaps someone else can answer the question.
1/2
How many different 4 letter permutations are there for the letters in the word toolroom?
120 four letter permutations if you don't allow more than one 'o' in the four letter
arrangement.
209 four letter permutations if you allow two, three and all four 'o'.
1.- Let set A = {t,l,r,m,}, and set B = {o,o,o,o}.
2.- From set A, the number of 4 letter permutations is 4P4 = 24.
3.- 3 letters from set A give 4P3 = 24, and one 'o' can take 4 different positions in the
word. That gives us 24x4 = 96 four letter permutations.
4.- In total, 24 + 96 = 120 different four letter permutations.
5.- If the other three 'o' are allowed to play, then you have 2 letters from set A that
give 4P2 = 12 permutations and two 'o' can take 4C2 = 6 position's, giving 12x6 = 72
four letter permutations.
6.- One letter from set A we have 4P1 = 4, each one can take 4 different positions, the
rest of the spaces taken by three 'o' gives 4x4 = 16 different permutations.
7.- The four 'o' make only one permutation.
8.- So now we get 72 + 16 + 1 = 89 more arrangements adding to a total of 89 + 120 = 209 different 4 letter arrangements made from the letters of the word toolroom.
[ nCr = n!/((n-r)!∙r!); nPr = n!/(n-r)! ]
a random sample
How many ways can you arrange the numbers 1468?
There are 4 digits. You can pick one of them as the first digit. So you can pick the first number 4 ways. Having picked the first number, the second number you can pick in only 3 ways. Having picked the first number and the second number you can pick the third number in only 2 ways. Having picked the first number and the second number and the third number you can pick the fourth number in only 1 way.
So total ways of arranging 1468 would be 4 x 3 x 2 x 1 = 24
What is the essence of finding the standard deviation of your data?
I am not entirely sure I understand correctly what you mean by "essence". However, the idea of finding the standard deviation is to determine, as a general tendency, whether most data points are close to the average, or whether there is a large spread in the data. The standard deviation means, more or less, "How far is the typical data point from the average?"
How do you get the probability of three dice rolls?
The probability of 3 specific dice rolls is the probability that each one will happen multiplied together. For instance, the probability of rolling 2 then 6 then 4 is the probability of all of these multiplied together: The probability of rolling 2 is 1/6. The probability of rolling 6 is 1/6. The probability of rolling 4 is 1/6. Multiply these together and we get the total probability as 1/216
What do you call eggnangle on a small spade?
The term "eggnangle" does not appear to be a widely recognized term in English. If you meant "egg angle," it may refer to the angle or position of an egg, possibly in a culinary or artistic context. However, if you are referring to a specific tool or concept involving a "small spade," please provide more context for a clearer answer.
What are the chances of getting tails at least twice if you throw a coin four times?
.5 or 50% probability (if not counting draws)
What is the P TTT on three flips of a coin?
P(T) = 1/2
Since you want to flip 3 T's in a row, you multiply.
P(TTT) = 1/2 * 1/2 * 1/2 = 1/8
Two coins are going to be tossed 780 times Predict the number of times 2 heads will be tossed?
Half the time they will be the same, half the time they will be different. Half of the time that they're the same they will be heads, half the time they are the same they will be tails. It's your homework, YOU figure it out. The way I figure it. There are four options:
1) heads / heads
2) heads / tails
3) tails / heads
4) tails / tails By process of chance, one out of four times both coins will be heads/heads. Therefore 780/4 = 195 times.
What are the percent of chance of rolling a six twice in a row on one six sided dice?
(1/6 * 1/6) * 100 = 2.8%
What is the the volume of a container that has 10 centimeters height and 10 centimeters width?
You need to know another dimension to find the volume. The volume is the length times height times width. You can only findthe area with two dimensions... the area would be 100cm2
See below for more info on area and volume.
What is the probability of a coin tossed four times to get 3 tails?
Using a calculator, plug numbers into the formula: nCk/2^n
4C3/2^4 = 4/16 = 1/4
Pascal's triangle can be used, if you are familiar with that (makes things much easier for a .5 probability scenario. See related link
Can a draw two card be placed on a draw four card in the game UNO?
No, if a draw 4 is placed, the player MUST draw 4, there is no way out of it. However if the player who placed the draw 4 has a card that is the current colour, you can challenge them, they then have to show there cards, if they do then they must draw 2, but if you are wrong you must draw 2 extra cards.